In a performance test, each of two cars takes 8.8 s to accelerate from rest to 26 m/s. Car A has a mass of 1442 kg, and car B has a mass of 1863 kg. Find the net average force that acts on (a) car A and (b) car B during the test.
1
Expert's answer
2012-10-11T09:21:48-0400
The acceleration of the both cars is

a = 26[m/s]/8.8[s] = 65/22 [m/s²].

According to the second Newton's law, the force acting on the first car is

F1 = 1442[kg] * 65/22 [m/s²] = 4260.5 N

and the force acting on the second car is

F2 = 1863[kg] * 65/22 [m/s²] = 5504.3 N.